What health club changes $44 sign-up fee and a $30 per month another health club charges are $99 sign-up fee and a $25 per month for what number of months is the cost the same?
Answer by ankor@dixie-net.com(22740)

What health club changes $44 sign-up fee and a $30 per month another health club charges are $99 sign-up fee and a $25 per month for what number of months is the cost the same?
:
let m = no. months for the cost to be the same
:
Equation for the cost of the 1st club
C(m) = 30m + 44
Equation for the cost of the 2nd
C(m) = 25m + 99
When will they cost the same?
30m + 44 = 25m + 99
30m - 25m = 99 - 44
5m = 55
m = 55/5
m = 11 months
:
:
Check this, find the actual cost of each
30(11) + 44 = $374
25(11) + 99 = $374